I have to simplify the following expression. Can you please tell me if I did this right?
sqrt45x^23 That is the square root of 45 x to the 23rd power.
Here is what I am getting. I'm not sure I am doing the "to the power" part right.
sqrt 9*x^16*5x^7
= sqrt9x^16 * sqrt5x^7
= 3x^4 sqrt5x^7
Am I right?

Answer by fractalier(2101)

Close. You pull out the LARGEST perfect square you can, noticing that any even power of x is a perfect square, so that
sqrt(45x^23) =
sqrt(9x^22)*sqrt(5x) =
3x^11 * sqrt(5x)
Also notice that the square root of a power is half the power, not the square root of the exponent...